OpenStudy (anonymous):

A step-up transformer has 300 turns on its primary coil and 90,000 turns on its secondary coil. If the current flowing through the secondary is .5 amps. What is the current flowing in the primary?

OpenStudy (michele_laino):

hint: here we have to apply this formula: \[\huge \frac{{{N_p}}}{{{N_s}}} = \frac{{{I_s}}}{{{I_p}}}\] where the subscript \(p\) stands for primary, and the subscript \(s\) stands for secondary. So we have: \[\Large {I_p} = \frac{{{N_s}}}{{{N_p}}}{I_s} = \frac{{90000}}{{300}} \cdot 0.5 = ...\;amperes\]